# An entanglement witness and an effective criterion for bipartite states of different dimensions

### v4

@@ 4.4 Example 4
$\gamma=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
@@
4.5 Example 5
This example is a continuation of the previous example and demonstrates that for the second PPT entangled state introduced in [23], the interval of entanglement (indicated as $10<b<$ 1), is indeed much larger and as in the previous example covers all positive and finite values of $b$ different from $1$. The state is given by
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{b}&.&.&1&.&.&.&.&.&.&.&.&.&.&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&b&.&.&.&.&.&.&.&.&1&.&.&.\\ \hline\cr.&1&.&.&b&.&.&.&.&.&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&\frac{1}{b}&.&.&1&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&1&.&.&b&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&1&.\\ \hline\cr.&.&.&1&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&.&.&.&.&.&.&.&.&.&.&1&.&.&b&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$
where $N=4b+\frac{4}{b}+8$, and its corresponding matrix is
$R=\frac{1}{(1+b)^{2}}\left(\frac{b}{2}I_{6}\oplus 0_{6}\oplus A\right),$
where $I_{6}$ and $0_{6}$ are respectively the identity matrix and the zero matrix in $C_{6}$ and
$A=\left(\begin{array}[]{ccc}-\frac{1}{8}(b-1)^{2}&-\frac{b^{2}+2b-3}{8\sqrt{3}}&-\frac{(b-1)b}{2\sqrt{6}}\\ \frac{3b^{2}-2b-1}{8\sqrt{3}}&-\frac{1}{24}(b-1)^{2}&-\frac{b^{2}+b-2}{6\sqrt{2}}\\ \frac{b-1}{2\sqrt{6}}&\frac{2b^{2}-b-1}{6\sqrt{2}}&-\frac{1}{12}(b-1)^{2}\\ \end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|b-1\right|}{4\left|b+1\right|}\ (2),\ \frac{\left|b\right|}{2(b+1)^{2}}\ (6),\ \frac{(b-1)^{2}}{4(b+1)^{2}}\ (1)\Big{\}},$
which is the same as the previous example and gives the same result.
@@ 4.6 Example 6
$\gamma=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
@@ 5 Conclusion
In this paper, a class of Entanglement Witnesses for two particle systems with different dimensions was presented. Specifically we showed in section (3), how the use of this class of EW’s can be reduced to a simple criterion, (equation (20)), which can be checked on the form of the density matrix. We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought.

### v5

@@ 5.2 Example 2
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
@@ 5.3 Example 3
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
@@ 6 Conclusion
Mehta’s lemma is a simple sufficient condition for a matrix to be positive. We have used this lemma to construct a large family of positive maps and characterized them in terms of their corresponding affine transformations, i.e. transformation of the so-called Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, equation (12). Using these positive maps, and by using the one-to-one correspondence between entanglement witnesses and the Choi matrix of positive maps, a multi-parameter family of Entanglement Witnesses for two particle systems with different dimensions was constructed. Specifically we showed how the use of this class of EW’s can be reduced to a simple criterion, (equation (35)), which can be checked on the form of the density matrix. We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought. We also rigorously proved that unital entanglement witnesses, i.e. those corresponding to unital positive maps, can detect a larger set of entangled states.
Acknowledgement: This research was supported in part by Iran National Science Foundation, under Grant No.4022322. The authors wish to thank S. Roofeh, A. Farmanian, and A. Tangestani, for many inspiring discussions during this project. The authors also thank Dariusz Chruscinski for his valuable comments.